### v1

# Retrieval Pivot Attacks in Hybrid RAG: Measuring and Mitigating Amplified Leakage from Vector Seeds to Graph Expansion

Abstract
Hybrid Retrieval-Augmented Generation (RAG) pipelines increasingly combine vector similarity search with knowledge graph expansion to support multi-hop reasoning. We show this composition introduces a distinct security failure mode: a semantically retrieved “seed” chunk can pivot via entity linking into sensitive graph neighborhoods, causing data leakage that does not exist in vector-only retrieval. We formalize this risk as Retrieval Pivot Risk (RPR) and define companion metrics Leakage@k, Amplification Factor (AF), and Pivot Depth (PD) to quantify leakage probability, magnitude, amplification over vector baselines, and structural distance to the first unauthorized node. We further present a taxonomy of four Retrieval Pivot Attacks that exploit the vector-to-graph boundary with injection budgets as small as 10–20 chunks, and we demonstrate that adversarial injection is not required: naturally shared entities (e.g., vendors, infrastructure, compliance standards) create cross-tenant pivot paths organically. In a synthetic multi-tenant enterprise corpus (1,000 documents; 2,785 graph nodes; 15,514 edges) evaluated with 500 queries, the undefended hybrid pipeline exhibits $\text{RPR}\approx 0.95$ and $\text{AF}(\epsilon)\approx 160\text{--}194\times$ relative to vector-only retrieval, with leakage occurring at $\text{PD}=2$ hops as a structural consequence of the bipartite chunk–entity topology. We validate these findings on the Enron email corpus (50,000 emails; 376,000 graph nodes; 2.3M edges), where undefended hybrid retrieval produces $\text{RPR}=0.695$ on benign queries—lower than the synthetic corpus due to sparser entity connectivity, but still affecting 70% of queries with a mean of 7.1 leaked items each. We propose five layered defenses and find that a single placement fix—per-hop authorization at the graph expansion boundary—eliminates all measured leakage ($\text{RPR}\to 0.0$) across both corpora, all queries, and all attack variants with negligible latency overhead, indicating the vulnerability is primarily a boundary enforcement problem rather than a defense complexity problem.
I Introduction
Enterprise adoption of Retrieval-Augmented Generation (RAG) has accelerated rapidly: 30–60% of enterprise AI use cases now rely on RAG architectures [9], and the vector database market reached \$1.73 billion in 2024 [14]. To improve multi-hop reasoning over complex organizational knowledge, practitioners increasingly deploy hybrid RAG pipelines that combine vector similarity search with knowledge graph expansion [16, 15]. These systems retrieve an initial set of text chunks via embedding similarity, then pivot through entity mentions into a knowledge graph to gather structurally related context before passing the assembled results to a large language model (LLM). We call the transition between vector retrieval and graph expansion the pivot boundary: the architectural point where a retrieved seed becomes a graph traversal capability.
This boundary is a boundary placement bug. The vector store enforces tenant policy before retrieval. The graph expansion must enforce it again after entity linking—otherwise an authorized seed chunk becomes an uncontrolled entry point into the knowledge graph. Consider an analyst at an engineering firm who queries about Kubernetes cluster configurations. Vector retrieval returns authorized engineering documents. But those documents mention shared entities (“CloudCorp,” “auth-service”) that also appear in the knowledge graph connected to confidential HR salary records, restricted security credentials, and financial audit data belonging to other tenants. A 2-hop graph expansion traverses through these shared entities into unauthorized neighborhoods, silently injecting sensitive cross-tenant data into the analyst’s context window.
Prior work falls into two separate buckets that do not address this boundary. (1) Vector-only attacks: PoisonedRAG achieves 90% attack success rate by injecting 5 malicious texts into vector stores [28]; CorruptRAG [24] and CtrlRAG [18] extend these results. (2) Graph-only attacks: GRAGPoison demonstrates 98% success via relation-centric poisoning within GraphRAG [8]; TKPA [20] and RAG-Safety [25] target graph-side integrity. (3) The hybrid boundary—how vector outputs become graph seeds—has not been systematically measured or formalized. The OWASP LLM Top 10 (2025) introduced LLM08 (Vector and Embedding Weaknesses) but does not address graph components [13]. MITRE ATLAS treats RAG as a monolithic system [11]. The SoK on RAG privacy explicitly identifies hybrid RAG security as an open problem [2].
This paper makes three contributions:
  1. 1.
    Composition vulnerability + metrics. We formalize Retrieval Pivot Risk (RPR) and companion metrics (AF, PD, Leakage@k, severity-weighted leakage) that quantify a compound attack surface emerging from composing two individually secure retrieval modalities. In our bipartite chunk–entity graph, all leakage occurs at PD == 2 hops—a structural signature of the pivot boundary (§V).
  2. 2.
    Cross-dataset validation + organic leakage. We demonstrate the vulnerability on both a synthetic enterprise corpus (1,000 documents, 4 tenants) and the Enron email corpus (50,000 emails, 5 departments), showing RPR == 0.95 and RPR == 0.70 respectively without adversarial injection. We present four Retrieval Pivot Attacks (A1–A4) exploiting the pivot boundary with injection budgets of 10–20 chunks (§IV, §VII).
  3. 3.
    Defense placement analysis. We evaluate five layered defenses (D1–D5) and show that authorization must be re-checked at the graph expansion boundary. D1 (per-hop authorization) alone eliminates all measured leakage ($\text{RPR}\to 0.0$) across both corpora, all queries, all attack variants, and metadata mislabel rates up to 5%. D2–D5 serve as utility optimizers and defense-in-depth layers (§VIII, §VII).
The complete codebase, data generators, attack implementations, defense suite, and experimental results are available at https://github.com/scthornton/hybrid-rag-pivot-attacks.
II Background
II-A Vector Retrieval in RAG
Standard RAG systems encode documents as dense vector embeddings using models such as all-MiniLM-L6-v2 [26] and retrieve the top-$k$ chunks by cosine similarity to the query embedding. In multi-tenant deployments, vector stores apply metadata prefilters (e.g., tenant ID) before similarity search, ensuring that retrieval respects organizational boundaries. This prefiltering makes vector-only RAG robust against cross-tenant leakage: our experiments confirm $\text{RPR}=0.0$ for vector-only pipelines across all evaluation queries on both synthetic and Enron corpora (§VII).
II-B Knowledge Graph RAG
GraphRAG [15] and related approaches construct knowledge graphs from document corpora via named entity recognition (NER) and relation extraction, then leverage graph structure for multi-hop reasoning. Systems like AgCyRAG [7] use LLM-driven graph traversal for complex queries. Graph-based retrieval provides superior comprehensiveness on multi-hop queries—86% versus 57% for vector RAG [16]—but introduces graph-specific attack surfaces: GRAGPoison [8] achieves 98% attack success through relation-centric poisoning, and TKPA [20] drops QA accuracy from 95% to 50% by modifying just 0.06% of corpus text.
II-C Hybrid RAG and the Pivot Boundary
Hybrid RAG combines both retrieval modalities in a pipeline:
  1. 1.
    Vector retrieval: Query embedding $\rightarrow$ cosine similarity $\rightarrow$ top-$k$ seed chunks $S(q)$.
  2. 2.
    Entity linking: Extract named entities from $S(q)$ via NER; map entity mentions to knowledge graph node IDs.
  3. 3.
    Graph expansion: BFS/DFS traversal from linked entity nodes to depth $d$, gathering structurally related nodes.
  4. 4.
    Context merge: Combine vector-retrieved chunks with graph-expanded context for LLM generation.
We define the pivot boundary as the transition between steps 1–2 and steps 3–4: the point where vector retrieval results become graph traversal seeds. This boundary is the attack surface we study. Vector-side prefilters operate before the pivot; graph expansion operates after it. If graph expansion does not independently enforce access controls, any entity mentioned in an authorized seed chunk becomes an uncontrolled entry point into the knowledge graph.
Running example. Consider an INTERNAL-clearance engineer at Acme Corp who queries “What infrastructure does auth-service depend on?” The vector store returns authorized Acme engineering documents mentioning auth-service. Entity linking maps “auth-service” to a shared entity node in the knowledge graph. BFS expansion then walks: auth-service $\to$ CloudCorp (shared vendor entity, hop 1) $\to$ Umbrella Security’s CONFIDENTIAL incident-response documents (hop 2). The engineer’s authorized query silently pulls CONFIDENTIAL cross-tenant data into the context—a 2-hop pivot through a legitimate shared entity.
III Threat Model
III-A System Model
We consider a multi-tenant enterprise hybrid RAG system with the following components:
  • A vector store (ChromaDB) containing document chunk embeddings with tenant and sensitivity metadata.
  • A knowledge graph (Neo4j) containing entity nodes, chunk nodes, and typed edges (MENTIONS, DEPENDS_ON, BELONGS_TO, RELATED_TO) extracted from the document corpus.
  • An entity linker that maps NER-extracted entities from vector-retrieved chunks to graph node IDs.
  • A graph expander that performs BFS traversal from linked entities to depth $d$ (default $d=2$), implemented via Neo4j’s apoc.path.spanningTree for hop-distance tracking.
  • Four organizational tenants with distinct data ownership boundaries.
  • Four sensitivity tiers: PUBLIC ($\approx$40% of corpus), INTERNAL (30%), CONFIDENTIAL (20%), RESTRICTED (10%).
III-B Attacker Capabilities
We consider two attacker models:
Injection attacker. The attacker holds a legitimate account in one tenant (acme_engineering) and can inject documents into the shared corpus through standard channels—wiki edits, ticket updates, shared document repositories. Injected documents undergo normal ingestion: chunking, NER-based entity extraction, embedding, and indexing into both vector and graph stores. The attacker’s injection budget is modest: 10–20 chunks, consistent with the budgets shown effective in prior work [28, 24].
No-injection attacker. The attacker holds a legitimate account and crafts queries that mention bridge entities (shared infrastructure names, vendor references, cross-team personnel) to maximize pivot probability through naturally occurring cross-tenant entity connections. This attacker requires zero document injection—the organic structure of the multi-tenant knowledge graph provides the pivot paths. Our evaluation shows that this attacker model achieves $\text{RPR}=0.95$ through benign queries alone (§VII), demonstrating that the vulnerability is structural rather than injection-dependent.
III-C Attacker Goals
  • G1: Cross-tenant access. Cause the retrieval context for a query in the attacker’s tenant to include chunks or entities belonging to other tenants.
  • G2: Sensitivity escalation. Cause an INTERNAL-clearance user’s context to include CONFIDENTIAL or RESTRICTED items.
  • G3: Amplified leakage. Achieve leakage in the hybrid pipeline that exceeds what is possible with vector-only retrieval (i.e., $\text{AF}>1$).
III-D Adaptive Attacker Considerations
An adaptive attacker aware of D1 (per-hop authorization) might attempt to bypass it through: (a) metadata spoofing—forging tenant or sensitivity labels during document injection, or (b) same-tenant escalation—crafting documents within the attacker’s own tenant that link to higher-sensitivity content within the same tenant boundary. D1 is robust against (a) if and only if the ingestion pipeline enforces metadata integrity (the metadata is assigned by the system, not the uploader). We evaluate D1’s robustness under metadata corruption in §VII-K and discuss mitigation strategies in §VIII-B. Same-tenant escalation (b) targets sensitivity boundaries rather than tenant boundaries and requires the attacker to already have write access to the target tenant’s corpus.
III-E Metadata Integrity Assumption
D1 (per-hop authorization) relies on node metadata—tenant labels and sensitivity tiers—being trustworthy. We assume metadata is system-assigned during document ingestion (the uploader cannot choose tenant or sensitivity labels). If an attacker can forge metadata, D1 can be bypassed. We evaluate robustness under random metadata corruption in §VII-K and discuss additional integrity mechanisms in §VIII-B.
III-F Scope
We measure unauthorized items in the retrieval context window, not exfiltrated tokens or generated text. We focus exclusively on the retrieval layer. We do not evaluate LLM generation quality, jailbreaking, or prompt injection. The attacker does not have direct database access, cannot modify model weights, and cannot alter the pipeline configuration. We omit a graph-only baseline (P2) because graph-only RAG without vector seeding is a fundamentally different retrieval paradigm—our question is amplification created by composition, not graph-only security, which is addressed by prior work [8, 20].
IV Retrieval Pivot Attacks
We define four non-adaptive attacks (A1–A4) that exploit the pivot boundary with increasing structural sophistication, and three adaptive attacks (A5–A7) that target the defense mechanisms themselves. Each attack operates within the threat model of §III: the attacker injects a small number of crafted documents that undergo standard ingestion.
IV-A A1: Seed Steering
Objective. Maximize the probability that attacker-crafted chunks are retrieved by target queries while embedding entity mentions that link to sensitive graph neighborhoods.
Mechanism. The attacker estimates the query centroid for target query families (e.g., “infrastructure monitoring”) and crafts 10 chunks with high semantic overlap to this centroid. Each chunk embeds 2 pivot entities—entities that are within 1–2 hops of sensitive nodes in the knowledge graph (e.g., k8s-prod-cluster, auth-service). Chunks are labeled PUBLIC with low provenance scores (0.3) to bypass any trust-based filtering. When retrieved, entity linking maps the embedded entity mentions to graph nodes, seeding expansion into sensitive neighborhoods.
Entry point. Vector retrieval (cosine similarity).
Pivot mechanism. Entity linking from retrieved chunk to graph node.
IV-B A2: Entity Anchor Injection
Objective. Force entity linking to create dense connections between injected chunks and sensitive graph neighborhoods.
Mechanism. The attacker identifies anchor entities adjacent to sensitive nodes (e.g., entities 1 hop from restricted credentials) and crafts chunks with dense entity mentions: 3+ mentions of the primary target entity per chunk, plus 2 related entities. NER extraction creates MENTIONS edges from each chunk to the target entities. Any query that retrieves one of these chunks—even incidentally—triggers graph expansion toward the sensitive area.
Entry point. Entity extraction during ingestion.
Pivot mechanism. MENTIONS edges from chunk to target entity nodes.
IV-C A3: Neighborhood Flooding
Objective. Inflate the local density around a target entity to increase the volume of sensitive content reachable through BFS expansion.
Mechanism. The attacker injects 20 chunks, each mentioning a high-value target entity near a sensitive subgraph. Each chunk also mentions a different neighbor entity, creating a dense web of edges. The result is a “graph gravity well”: the target entity accumulates many more edges than the graph average, causing BFS expansion to gather a larger neighborhood when traversing through this entity. Note that our pipeline uses uniform BFS traversal (not degree-weighted or PageRank-based expansion), so the flooding increases reachable volume rather than biasing traversal order.
Entry point. Graph topology manipulation via ingestion.
Pivot mechanism. Inflated local density increases BFS expansion volume.
IV-D A4: Bridge Node Attack
Objective. Create artificial cross-tenant edges that enable graph traversal from the attacker’s tenant into a target tenant.
Mechanism. The attacker crafts 15 chunks that co-mention entities from both the attacker’s tenant and the target tenant within the same document. NER extraction creates entity nodes on both sides of the tenant boundary, and relation extraction creates RELATED_TO edges between them. After ingestion, BFS expansion from the attacker’s authorized subgraph can traverse these artificial bridge edges into the target tenant’s data.
Entry point. Cross-tenant entity co-mention.
Pivot mechanism. Artificial RELATED_TO edges spanning tenant boundaries.
Algorithm 1 formalizes the general pivot attack framework.
Algorithm 1 General Retrieval Pivot Attack
1:Target query family $\mathcal{Q}_{T}$, injection budget $B$, pivot entities $E_{p}$, target neighborhood $N_{T}$
2:Injected chunks $C_{\text{atk}}$
3:for $i=1$ to $B$ do
4:  $c_{i}\leftarrow$ CraftChunk($\mathcal{Q}_{T}$, $E_{p}$) $\triangleright$ High similarity to $\mathcal{Q}_{T}$
5:  $c_{i}.\text{entities}\leftarrow$ EmbedEntities($E_{p}$, $N_{T}$) $\triangleright$ Mentions near $N_{T}$
6:  $c_{i}.\text{sensitivity}\leftarrow$ PUBLIC
7:  $c_{i}.\text{provenance}\leftarrow 0.3$
8:  $C_{\text{atk}}\leftarrow C_{\text{atk}}\cup\{c_{i}\}$
9:end for
10:Ingest($C_{\text{atk}}$) $\triangleright$ Standard pipeline: chunk, NER, embed, index
11:return $C_{\text{atk}}$
V Metrics
We introduce metrics to quantify retrieval pivot risk. Let $q$ denote a query, $u$ a user with tenant $t_{u}$ and clearance level $\ell_{u}$, $S_{k}(q)$ the top-$k$ context set, and $\text{Sensitive}(x,u)$ a predicate that is true when item $x$ has sensitivity $\ell_{u}$ or tenant $\neq t_{u}$ (entity nodes with empty-string tenants are excluded from cross-tenant counts, as they are tenant-neutral shared concepts—not leaked items themselves but rather the pivot bridges through which leakage propagates). Let $\text{Seeds}(q)\subseteq S_{k}(q)$ denote the set of entity nodes that are linked from vector-retrieved chunks via NER extraction—these are the graph entry points that initiate BFS expansion.
Item serialization. Each item in $S_{k}(q)$ is a node serialized for the LLM context: chunk nodes are serialized as their source text; entity nodes are serialized as {name, type, top-3 relations}. Context size counts the total number of serialized nodes (chunks + entities) presented to the LLM. The default pipeline configuration (depth $d=2$, branching $\leq 10$, total $\leq 100$ nodes, top-k=10k{=}10 vector results) produces $\sim$}110 items per query in the undefended hybrid pipeline (P3). Table I summarizes the notation used throughout.
TABLE I: Notation summary.
Symbol Definition
$q$ Query
$u$ User with tenant $t_{u}$ and clearance $\ell_{u}$
$S_{k}(q)$ Top-$k$ context set (serialized items)
$\text{Seeds}(q)$ Entity nodes linked from vector-retrieved chunks
$\text{Sensitive}(x,u)$ $x$ has $\ell_{x}>\ell_{u}$ or $t_{x}\neq t_{u}$
RPR Prob. of any sensitive item in context
Leakage@k Count of sensitive items in context
SWL Severity-weighted leakage
AF($\epsilon$) Leakage ratio: hybrid / max(vector, $\epsilon$)
PD Min hops from seed to first sensitive node
V-A Retrieval Pivot Risk (RPR)
RPR measures the probability that a query’s retrieval context contains any unauthorized item:
$\text{RPR}(u)=\Pr_{q\sim\mathcal{Q}}\left[\exists x\in S_{k}(q):\text{Sensitive}(x,u)\right]$ (1)
RPR is operationalized as the fraction of evaluation queries whose context contains at least one sensitive item. We report RPR with 95% bootstrap confidence intervals (10,000 resamples).
V-B Leakage@k
Leakage@k counts the number of unauthorized items in the context:
$\text{Leakage@k}(q,u)=\left|\{x\in S_{k}(q):\text{Sensitive}(x,u)\}\right|$ (2)
While RPR captures whether leakage occurs, Leakage@k captures its severity. A context with 1 leaked item and one with 20 both yield $\text{RPR}=1$, but their Leakage@k values differ by $20\times$.
V-C Severity-Weighted Leakage
Not all leakage is equally harmful. A PUBLIC-clearance user receiving INTERNAL content is a lesser violation than receiving RESTRICTED content. We define severity-weighted leakage as:
$\text{SWL}(q,u)=\sum_{\begin{subarray}{c}x\in S_{k}(q)\\ \text{Sensitive}(x,u)\end{subarray}}w(x,u)$ (3)
where the per-item weight $w(x,u)$ captures two distinct policy violations:
$w(x,u)=\begin{cases}\ell_{x}-\ell_{u}&\text{if }\ell_{x}>\ell_{u}\\ 1&\text{if }t_{x}\neq t_{u}\wedge\ell_{x}\leq\ell_{u}\end{cases}$ (4)
Here $\ell_{x}$ is the sensitivity level (PUBLIC=0, INTERNAL=1, CONFIDENTIAL=2, RESTRICTED=3) and $t_{x}$ the tenant of item $x$. Over-clearance items contribute weight proportional to the sensitivity gap. Cross-tenant items that do not exceed the user’s clearance receive a penalty of 1, reflecting the policy violation of accessing another tenant’s data regardless of sensitivity. In our experiments, the INTERNAL-clearance user observing RESTRICTED items produces weight 2 per item, CONFIDENTIAL produces weight 1, and cross-tenant PUBLIC items produce weight 1.
V-D Amplification Factor
AF quantifies the leakage increase that hybrid retrieval introduces compared to vector-only retrieval:
$\text{AF}=\frac{\mathbb{E}[\text{Leakage@k}]_{\text{hybrid}}}{\mathbb{E}[\text{Leakage@k}]_{\text{vector}}}$ (5)
When the vector baseline produces zero leakage (which our experiments confirm), $\text{AF}=\infty$ for any nonzero hybrid leakage. To provide a finite, plottable alternative, we also report $\text{AF}(\epsilon)$:
$\text{AF}(\epsilon)=\frac{\mathbb{E}[\text{Leakage@k}]_{\text{hybrid}}}{\max(\mathbb{E}[\text{Leakage@k}]_{\text{vector}},\;\epsilon)}$ (6)
with $\epsilon=0.1$, alongside the absolute difference $\Delta\text{Leakage}=\mathbb{E}[\text{Leakage@k}]_{\text{hybrid}}-\mathbb{E}[\text{Leakage@k}]_{\text{vector}}$.
[…]
V-E Pivot Depth (PD)
PD measures the minimum graph distance from a seed node to the first sensitive node reached during expansion:
$\text{PD}(q)=\min\{d(s,x):s\in\text{Seeds}(q),\;x\in S_{k}(q),\;\text{Sensitive}(x,u)\}$ (7)
where $d(s,x)$ is the shortest-path distance in the knowledge graph from seed $s$ to node $x$, computed via apoc.path.spanningTree which yields paths with explicit hop counts. We report PD as a distribution (min, median, max) across queries that exhibit leakage, rather than a single summary statistic. Operationally, PD identifies the minimum traversal depth at which enforcement must occur: if all leakage has $\text{PD}=d$, limiting expansion to depth <d<d or inserting an authorization check at depth $d$ is sufficient to prevent it.
VI Experimental Setup
VI-A Synthetic Enterprise Corpus
We generate a multi-tenant enterprise corpus of 1,000 documents across four organizational tenants: acme_engineering, globex_finance, initech_hr, and umbrella_security. Documents are produced by 12 domain-specific generators (3 per tenant) that embed realistic entity mentions—system names, personnel, projects, compliance standards—using curated entity pools.
Sensitivity tiers follow a realistic distribution: PUBLIC (40%), INTERNAL (30%), CONFIDENTIAL (20%), and RESTRICTED (10%). Each document includes ground-truth entity annotations and sensitivity labels.
Bridge entities. We inject 15 bridge entities across 5 categories (shared vendors, shared infrastructure, shared personnel, shared compliance standards, and shared projects) that naturally appear in documents across multiple tenants. For example, “CloudCorp” appears in both engineering and finance documents, and “auth-service” appears in both engineering and security documents. After NER-based entity extraction, the knowledge graph contains 40 naturally shared entities across tenant boundaries—not through adversarial injection, but through legitimate cross-team references that any multi-tenant organization would exhibit.
VI-B Knowledge Graph Construction
Documents undergo chunking (300-token windows, 50-token overlap), producing $\sim$}2,000 chunks from 1,000 source documents (average $\sim$}2 chunks per document). Chunks are processed through spaCy NER extraction (en_core_web_sm), two-pass relation extraction (ground-truth relation resolution followed by pattern-based extraction across 5 typed relations: DEPENDS_ON, OWNED_BY, BELONGS_TO, CONTAINS, DERIVED_FROM, plus RELATED_TO fallback), and embedding via all-MiniLM-L6-v2 (384 dimensions). The resulting knowledge graph contains 2,785 nodes (785 entity nodes + 2,000 chunk nodes) and 15,514 edges (7,386 extracted relations plus MENTIONS, CONTAINS, and BELONGS_TO structural edges added during graph construction).
VI-C Enron Email Corpus
To validate that Retrieval Pivot Risk is not an artifact of synthetic construction, we evaluate on the Enron email corpus—a public-record dataset of $\sim$}500,000 corporate emails from $\sim$}150 Enron employees, released during the 2001 FERC investigation. We subsample to 50,000 emails from the most active employees and assign tenants based on departmental structure:
  • trading — Trading, West/East Power Trading
  • legal — Legal, Government Affairs
  • finance — Finance, Risk Management, Accounting
  • energy_services — Energy Services, Pipeline, ENA
  • executive — Executive, Office of the Chairman
Sensitivity labels are assigned by keyword matching: RESTRICTED (attorney-client privilege markers, password shares, strategy memos), CONFIDENTIAL (deal negotiations, valuations, board communications), INTERNAL (standard departmental communications), PUBLIC (company-wide announcements). The resulting distribution (92% PUBLIC, 6.7% CONFIDENTIAL, 0.7% RESTRICTED, 0.3% INTERNAL) is skewed compared to the synthetic corpus—reflecting realistic email traffic where most messages are routine.
Chunking (300-token windows, 50-token overlap) produces 152,064 chunks. spaCy NER extraction (en_core_web_sm) identifies 2.07M entity mentions. The resulting knowledge graph contains 376,000 nodes (174,000 entity nodes + 152,000 chunk nodes + 50,000 document nodes) and 2.3M edges. Nineteen entities are naturally shared across departmental boundaries—cross-department executives (Ken Lay, Jeff Skilling, Andy Fastow), external organizations (Arthur Andersen, Vinson & Elkins), deal names (Project Raptor, LJM2), and internal systems (EnronOnline).
Key differences from synthetic. The Enron graph is $135\times$ larger than the synthetic graph (376K vs. 2.8K nodes) but has fewer intentional bridge entities (19 vs. 40). Email text produces noisier NER extractions (dates, monetary amounts, partial names) than curated synthetic documents, creating more spurious entity connections but also more entity fragmentation. These properties make the Enron corpus a stress test for whether the pivot vulnerability generalizes beyond controlled conditions.
VI-D Pipeline Variants
We evaluate 7 pipeline configurations (Table II):
TABLE II: Pipeline variants and their defense configurations.
ID Pipeline Defenses
P1 Vector-only Tenant prefilter
P3 Hybrid baseline None
P4 Hybrid + D1 Per-hop authz
P5 Hybrid + D1,D2 + Edge allowlist
P6 Hybrid + D1–D3 + Budgeted traversal
P7 Hybrid + D1–D4 + Trust weighting
P8 Hybrid + D1–D5 + Merge filter
We omit a graph-only baseline (P2) because graph-only retrieval without vector seeding is a fundamentally different paradigm that does not involve the pivot boundary under study. Its security properties are addressed by prior work on GraphRAG attacks [8, 20].
VI-E Evaluation Protocol
Synthetic corpus. We evaluate each pipeline on 500 template-generated queries: 350 benign queries (standard domain questions stratified across 4 tenants and 3 clearance levels) and 150 adversarial queries (queries that mention bridge entities or target cross-tenant pivot paths, stratified across attack types A1–A4). The evaluation user belongs to acme_engineering with INTERNAL clearance.
Enron corpus. We evaluate P1, P3, and P4 on 200 queries (100 benign + 100 adversarial) generated from department-specific templates. The evaluation user belongs to trading with INTERNAL clearance.
[…]
Statistical methodology. All metrics are reported with 95% bootstrap confidence intervals (10,000 resamples, percentile method, seed 42). We use non-parametric bootstrap because RPR values near 0 or 1 violate normal approximations, and leakage distributions are highly skewed. Our query sets (500 synthetic, 200 Enron) each provide power >0.80>0.80 for detecting 5pp differences (Appendix F provides sample size justification, multiple comparison corrections, and $\epsilon$ sensitivity analysis).
Reproducibility. All experiments use a single fixed random seed (42) for corpus generation, query sampling, and bootstrap resampling. Pipeline configurations are built programmatically (not from YAML) to ensure parameter consistency across variants. Appendix A provides the full reproduction sequence; Appendix B details the corpus generator; Appendix C lists all query templates. The codebase includes 255 passing unit tests.
VI-F Attack Evaluation
We evaluate all four attacks (A1–A4) against the undefended hybrid pipeline (P3) and three defense configurations (P4, P6, P8). For each attack, payloads are injected into a clean corpus, and 10 adversarial queries are executed against each pipeline variant. After each attack evaluation, the graph is rebuilt from the clean corpus to prevent cross-contamination between attack experiments.
VII Results
VII-A Hybrid RAG Amplifies Leakage
Table III presents security metrics across all pipeline variants. The central finding is stark: the undefended hybrid pipeline (P3) leaks massively while the vector-only baseline (P1) leaks nothing.
P1 achieves $\text{RPR}=0.0$ on both query sets—the vector store’s tenant prefilter prevents any cross-tenant or sensitivity-escalated content from reaching the context. In contrast, P3 achieves $\text{RPR}=0.954$ [95% CI: 0.931, 0.974] on 350 benign queries (334 of 350 queries produce leakage) and $\text{RPR}=0.947$ [0.907, 0.980] on 150 adversarial queries (142 of 150 queries leak). Mean Leakage@k reaches 16.0 items for benign queries and 19.4 for adversarial queries—meaning that roughly 15–18% of the 110-item context consists of unauthorized content. Severity-weighted leakage averages 22.9 (benign) and 26.4 (adversarial), indicating that leaked items skew toward higher sensitivity tiers.
TABLE III: Defense ablation: security metrics across pipeline variants (500 queries: 350 benign + 150 adversarial). RPR = Retrieval Pivot Risk with 95% bootstrap CI, Leak = mean Leakage@k, SWL = severity-weighted leakage, PD = Pivot Depth (hops), Ctx = mean context size. “–” indicates no leakage occurred (PD undefined).
Benign Queries ($n=350$) Adversarial Queries ($n=150$)
Variant RPR [CI] Leak SWL PD Ctx RPR [CI] Leak SWL PD Ctx
P1 (Vector) 0.000 0.0 0.0 10 0.000 0.0 0.0 10
P3 (Hybrid) .954 [.931,.974] 16.0 22.9 2.0 110 .947 [.907,.980] 19.4 26.4 2.0 110
P4 (+D1) 0.000 0.0 0.0 56 0.000 0.0 0.0 50
P5 (+D1,D2) 0.000 0.0 0.0 57 0.000 0.0 0.0 51
P6 (+D1–D3) 0.000 0.0 0.0 29 0.000 0.0 0.0 28
P7 (+D1–D4) 0.000 0.0 0.0 28 0.000 0.0 0.0 24
P8 (All) 0.000 0.0 0.0 20 0.000 0.0 0.0 20
[…]
Figure 1 shows RPR with bootstrap confidence intervals across pipeline variants.
Figure 1: Retrieval Pivot Risk with 95% bootstrap CIs across pipeline variants. P3 (undefended hybrid) shows $\text{RPR}\approx 0.95$. All defended variants (P4–P8) achieve $\text{RPR}=0.0$.
VII-B The PD = 2 Structural Signature
All leakage in P3 occurs at exactly $\text{PD}=2$ hops: the PD distribution is $(\min=2,\text{median}=2,\max=2)$ across all 476 leaking queries. This uniformity is not coincidental—it is a structural consequence of the bipartite chunk-entity graph topology in our construction. The pivot path is:
  1. 1.
    Hop 0: Authorized seed chunk (retrieved by vector similarity, passes tenant prefilter).
  2. 2.
    Hop 1: Shared entity node (linked via NER from seed chunk text; entity nodes carry no tenant ownership).
  3. 3.
    Hop 2: Unauthorized chunk (connected to the shared entity via MENTIONS edge; belongs to a different tenant or higher sensitivity tier).
This 2-hop pattern is inherent to any hybrid RAG system that constructs a bipartite graph between chunks and entities with expansion depth $d\geq 2$: the entity-linking step creates a structural bridge between vector-retrieved content and graph-stored content. The PD=2 finding is structural in our bipartite construction—knowledge graphs with richer entity-to-entity relationships (e.g., ontological hierarchies, multi-hop inference chains) may exhibit leakage at deeper pivot depths. We discuss how PD varies with traversal parameters in §VII-I.
VII-C Organic Leakage: No Injection Required
A critical observation from our 350 benign queries is that $\text{RPR}=0.954$ without any adversarial injection. The 40 naturally shared entities across tenant boundaries—shared infrastructure, vendors, personnel, compliance standards, and projects—provide sufficient pivot paths for massive leakage through ordinary queries. In our corpus, 334 of 350 benign queries (95.4%) trigger cross-tenant leakage through organic entity connections. This means the vulnerability is structural: it exists in any multi-tenant hybrid RAG deployment where tenants share real-world entities, regardless of whether an attacker injects content.
Bridge category analysis. To understand which types of shared entities drive leakage, we analyze the hop-1 pivot nodes in each leaking query’s traversal path and classify them against the 5 bridge entity categories (Table IV).
TABLE IV: Organic leakage by bridge entity category under P3 (no injection). Queries = leaking queries with that bridge type at hop 1. Leak = attributed leaked items.
Bridge Category Benign Adversarial
Queries Leak Queries Leak
Personnel 104 1,321 67 1,244
Compliance 46 782 33 489
Infrastructure 43 406 16 213
Vendor 0 0 30 218
Project 7 87 24 228
(No bridge at hop 1) 175 22
Total leaking 334 5,595 142 2,907
Personnel entities (shared employees like “Maria Chen”) dominate: they appear at hop 1 in 31% of benign leaking queries and 47% of adversarial leaking queries, accounting for 23.6% and 42.8% of attributed leakage respectively. Compliance entities (SOC2, PCI-DSS, ISO27001) rank second. Notably, 52% of benign leaking queries have no recognized bridge entity at hop 1—non-bridge entities (monetary amounts, dates, generic organizational terms extracted by spaCy NER) also create unintended cross-tenant paths, indicating that the leakage risk is broader than just named shared entities.
VII-D Connectivity Sensitivity
To measure how shared-entity density affects leakage, we regenerate the corpus with bridge entity counts $\in\{0,5,10,15,25,40\}$, rebuild all indexes for each count, and run 100 adversarial queries through P3 (Table V).
TABLE V: Connectivity sweep: effect of bridge entity count on P3 RPR and mean Leakage@k. RPR is remarkably stable; mean leakage increases monotonically with connectivity.
Bridges P3 RPR Mean Leak PD
0 0.93 21.5 2.0
5 0.95 26.6 2.0
10 0.95 30.5 2.0
15 0.95 31.5 2.0
25 0.95 32.5 2.0
40 0.94 34.5 2.0
Two findings emerge. First, RPR is remarkably stable (0.93–0.95) regardless of bridge count—even with zero intentional bridge entities, organic entity overlap from spaCy NER produces RPR =0.93=0.93. Second, mean leakage scales monotonically from 21.5 (0 bridges) to 34.5 (40 bridges), a 60% increase. Bridge entities do not enable leakage (BFS already reaches cross-tenant nodes through generic NER entities), but they amplify the volume of leaked content per query. PD remains uniformly 2.0 across all bridge counts, confirming the structural bipartite pivot.
VII-E Embedding Model Sensitivity
To verify that the pivot vulnerability is structural rather than embedding-dependent, we repeat the P1 and P3 evaluations using all-mpnet-base-v2 (768 dimensions, higher retrieval quality) in addition to our default all-MiniLM-L6-v2 (384 dimensions). We rebuild the ChromaDB collection for each model and run the full 500-query evaluation (Table VI).
TABLE VI: Embedding model sensitivity: P3 RPR and mean Leakage@k across two embedding models. The vulnerability persists regardless of embedding quality.
Model Benign Adversarial
RPR Leak RPR Leak
MiniLM-L6 (384d) 0.954 16.0 0.947 19.4
MPNet (768d) 0.994 19.4 0.980 28.4
The higher-quality MPNet model actually increases RPR: from 0.954 to 0.994 on benign queries, and from 0.947 to 0.980 on adversarial queries. Mean leakage also rises—from 16.0 to 19.4 (benign) and from 19.4 to 28.4 (adversarial). Better embeddings retrieve more relevant seed chunks, which in turn mention more entities, which seed more graph expansion into cross-tenant neighborhoods. Both models produce $\text{PD}=2.0$ uniformly, confirming that the bipartite pivot structure is model-independent. P1 achieves $\text{RPR}=0.0$ under both models, confirming that vector-side tenant prefiltering remains effective regardless of embedding quality.
VII-F Attack Evaluation
Table VII presents RPR under each attack across pipeline variants. All four attacks achieve $\text{RPR}=1.0$ against the undefended hybrid pipeline (P3)—every adversarial query produces cross-tenant leakage. Critically, all four attacks achieve $\text{RPR}=0.0$ against every defended variant (P4, P6, P8).
TABLE VII: RPR under each attack type across pipeline variants. Mean Leakage@k shown in parentheses. All attacks fail against D1-defended pipelines.
Attack P3 P4 P6 P8
A1 (Seed Steer) 1.00 (20.5) 0.00 (0.0) 0.00 (0.0) 0.00 (0.0)
A2 (Entity Anchor) 1.00 (20.5) 0.00 (0.0) 0.00 (0.0) 0.00 (0.0)
A3 (Nbhd Flood) 1.00 (20.5) 0.00 (0.0) 0.00 (0.0) 0.00 (0.0)
A4 (Bridge Node) 1.00 (20.5) 0.00 (0.0) 0.00 (0.0) 0.00 (0.0)
Table VIII provides injection details for each attack. The attacks span a range of injection budgets (9–20 chunks) and entity strategies (1–2 target entities, 1–6 MENTIONS edges created).
TABLE VIII: Attack injection details. All attacks achieve RPR=1.0 on P3 and RPR=0.0 on P4/P6/P8.
Attack Payloads Chunks Entities Mentions
A1 9 9 1 3
A2 10 10 1 6
A3 20 20 1 4
A4 15 15 2 6
Two aspects of this uniformity are themselves findings. Leakage is bounded by the expansion window, not the attack mechanism: all four attacks produce identical Leakage@k ($\approx$20.5) against P3 because the bottleneck is the unguarded graph traversal—all four simply steer queries into the same undefended 2-hop expansion path, and the total_nodes budget (100) caps the leaked volume. This means more sophisticated injection strategies yield no additional leakage beyond what organic entity overlap already provides. Second, D1 blocks all four strategies because it operates on node properties (tenant, sensitivity) rather than graph structure (paths, edges). The attacks manipulate paths—optimizing similarity (A1), creating dense connections (A2), inflating density (A3), or bridging boundaries (A4)—but D1 filters on properties at each hop, making the path irrelevant.
VII-G Adaptive Attacks (A5–A7)
We extend the attack taxonomy with three adaptive strategies that target the defense mechanisms themselves (Table IX).
TABLE IX: RPR under adaptive attacks (A5–A7) across pipeline variants. A5 tests metadata forgery at three rates; A6 tests entity manipulation; A7 tests query manipulation. Leakage@k shown in parentheses.
Attack Rate P3 P4 P7 P8
A5 1% 1.00 (7.8) 0.00 (0.0) 0.00 (0.0) 0.00 (0.0)
A5 5% 1.00 (7.8) 0.00 (0.0) 0.00 (0.0) 0.00 (0.0)
A5 10% 1.00 (7.8) 0.00 (0.0) 0.00 (0.0) 0.00 (0.0)
A6 1.00 (7.8) 0.00 (0.0) 0.00 (0.0) 0.00 (0.0)
A7 query-only (no injection)
A5 (Metadata Forgery) relabels injected nodes with the target tenant’s name to bypass D1’s tenant check. At forgery rates of 1%, 5%, and 10%, A5 achieves $\text{RPR}=1.0$ against P3 but $\text{RPR}=0.0$ against all defended pipelines (P4, P7, P8). D1 remains effective because the forged metadata only affects the attacker’s own injected nodes—the organic entity nodes that create the pivot path still carry empty-string tenant labels and are filtered.
A6 (Entity Manipulation) creates documents mentioning entities from the target tenant’s namespace, attempting to create new shared entity nodes. A6 also fails against D1: the newly created entity nodes still carry empty-string tenant labels, and the defense filters them regardless of how they were created.
A7 (Query Manipulation) crafts queries mentioning target tenant entity names to steer NER-based entity linking toward sensitive neighborhoods. This is a query-only attack (no injection) and produces the same RPR as organic benign queries—confirming that the vulnerability is in the graph expansion, not in query processing.
VII-H Generation Impact
To measure whether leaked context actually contaminates LLM-generated answers, we evaluate three production LLMs on 10 queries with known leakage (Table X).
TABLE X: Generation contamination metrics. ECR (Entity Contamination Rate) measures leaked entity names in answers. ILS (Information Leakage Score) measures semantic similarity to leaked content. FCR (Factual Contamination Rate) and GRR (Generation Refusal Rate) assess factual leakage and model self-censoring.
LLM ECR ILS FCR GRR Cost
GPT-4o 0.077 0.305 0.050 0.800 \$0.08
Claude Sonnet 4.5 0.307 0.346 0.065 0.100 \$0.18
DeepSeek-V3 0.124 0.336 0.190 0.200 \$0.01
Note: 10 queries with leakage evaluated per provider on synthetic dataset. Judge model: same provider as generator for each row. GPT-4o exhibits high refusal rate (GRR = 0.80), largely ignoring leaked context.
The results reveal model-dependent contamination behavior. Claude Sonnet 4.5 exhibits the highest Entity Contamination Rate ($\text{ECR}=0.31$) and lowest Generation Refusal Rate ($\text{GRR}=0.10$), indicating it readily incorporates leaked entities into generated answers. GPT-4o shows the opposite pattern: $\text{ECR}=0.08$ and $\text{GRR}=0.80$, largely ignoring leaked context. DeepSeek-V3 falls between ($\text{ECR}=0.12$, $\text{GRR}=0.20$) but produces the highest Factual Contamination Rate ($\text{FCR}=0.19$), meaning it synthesizes leaked facts into answers more than the other models. These findings confirm that retrieval-level leakage translates to generation-level contamination for most LLMs, and that the severity depends on the model’s tendency to incorporate all context versus selectively ignoring irrelevant content.
VII-I Traversal Parameter Sweep
To understand which traversal parameters control leakage, we sweep across 27 configurations: depth $\in\{1,2,3\}$, branching factor $\in\{5,10,25\}$, and total node budget $\in\{25,50,100\}$, running 100 adversarial queries per configuration against the undefended hybrid pipeline (P3). Figure 2 shows the results.
Figure 2: Traversal parameter sweep: context size vs. latency, colored by RPR. The total node budget is the primary leakage-controlling parameter.
Three findings emerge:
Total node budget is the primary control. Regardless of depth or branching, $\text{total\_nodes}=25$ yields $\text{RPR}=0$ (insufficient expansion to reach cross-tenant content), $\text{total\_nodes}=50$ yields $\text{RPR}\approx 0.66$, and $\text{total\_nodes}=100$ yields $\text{RPR}\approx 0.93$. This is because the total node budget caps how many graph nodes are gathered, and once this cap prevents expansion from reaching the 2-hop cross-tenant chunks, leakage is eliminated.
[…]
Branching factor is irrelevant given a total node cap. At fixed depth and total node budget, branching $\in\{5,10,25\}$ produces identical RPR values. The BFS expansion fills the node budget regardless of how many children are explored per node.
VII-J Latency and Overhead
Table XI presents latency and context size across pipeline variants.
TABLE XI: Latency (ms) and context size across pipeline variants.
Variant p50 p95 Mean Ctx
Benign Queries
P1 (Vector) 12.6 19.9 16.2 10
P3 (Hybrid) 26.7 38.1 32.4 110
P4 (+D1) 26.5 34.4 30.4 56
P6 (+D1–D3) 23.1 29.8 26.4 29
P8 (All) 22.7 28.6 25.6 20
Adversarial Queries
P1 (Vector) 12.3 34.6 23.5 10
P3 (Hybrid) 27.3 40.2 33.7 110
P4 (+D1) 26.3 32.2 29.3 50
P6 (+D1–D3) 23.1 33.7 28.4 28
P8 (All) 22.9 30.1 26.5 20
[…]
VII-K Metadata Integrity Stress Test
[…]
Figure 3: D1 robustness under metadata corruption. RPR remains 0.0 even at 5% mislabel rate. Context size decreases slightly as erroneously up-labeled nodes are filtered.
VII-L Defense Ablation
The defense stack shows a clear pattern (Figure 4):
  • D1 alone (P4): RPR drops from $0.95$ to $0.0$. Context reduces from 110 to 50–56 items (the unauthorized items are removed; authorized graph-expanded content is retained).
  • D1 + D2 (P5): No additional security improvement; context remains similar (51–57). Edge allowlisting provides defense in depth but does not further reduce leakage already eliminated by D1.
  • D1–D3 (P6): Context drops to 28–29. Budgeted traversal caps the number of expanded nodes, reducing context noise from authorized but irrelevant graph content.
  • D1–D4 (P7): Context drops to 24–28. Trust-weighted filtering removes low-provenance nodes.
  • D1–D5 (P8): Context reaches 19–20 items. The merge filter provides a final defense-in-depth check, reducing context by 82% from the undefended baseline.
Figure 4: Mean context size under progressive defenses. D1 alone reduces context from 110 to 50–56 items (removing unauthorized content). D3–D5 further reduce noise, reaching 19–20 items with the full defense stack.
The key insight is that D1 is both necessary and sufficient for security (eliminating all leakage), while D2–D5 serve as utility optimizers (reducing context noise and providing defense in depth).
VII-M Utility Impact
A critical question for practitioners is whether defenses degrade retrieval quality. Since our evaluation corpus uses synthetic documents without human-annotated relevance labels, we measure utility through two proxy metrics: (1) authorized context size—the number of items passing authorization checks in the final context, and (2) authorization rate—the fraction of context items that are authorized (i.e., not leaked). Together these quantify how much useful content each defense preserves and how much noise it eliminates.
Table XII presents the security-utility tradeoff. P3 (undefended hybrid) returns 110 items per query, but only 85.5% are authorized—the remaining 16 items (14.5%) are cross-tenant or over-clearance leakage. P1 (vector-only) returns 10 fully authorized items, establishing the vector-only utility baseline.
[…]
Variant RPR Ctx Auth. Auth. Rate Retention
P1 (vector) 0.000 10 10.0 1.00
P3 (hybrid) 0.954 110 94.0 0.85 1.00
P4 (+D1) 0.000 56 56.0 1.00 0.60
P6 (+D1–3) 0.000 29 29.0 1.00 0.31
P8 (+D1–5) 0.000 20 20.0 1.00 0.21
[…]
VII-N Cross-Dataset Validation
To verify that Retrieval Pivot Risk is not an artifact of synthetic corpus construction, we repeat the baseline evaluation (P1, P3, P4) on the Enron email corpus described in §VI-C. Table XIII compares results across both corpora.
TABLE XIII: Cross-dataset baseline: retrieval security metrics for undefended (P3) and defended (P4) hybrid RAG across synthetic and Enron corpora. RPR = Retrieval Pivot Risk, Leak = mean leaked items per query, PD = Pivot Depth (hops), Ctx = mean context size.
Dataset Pipeline RPR Leak PD Ctx
Synthetic P1 (Vector) 0.000 0.0 10
P3 (Hybrid) 0.954 16.0 2.0 110
P4 (+D1) 0.000 0.0 56
Enron P1 (Vector) 0.000 0.0 10
P3 (Hybrid) 0.695 7.1 2.0 45
P4 (+D1) 0.000 0.0 25
Note: Synthetic corpus: 1000 documents, 4 tenants, 40 bridge entities. Enron corpus: 50,000 emails, 5 departments, 19 cross-department entities. Both corpora evaluated with 200 benign queries.
Three findings emerge from the cross-dataset comparison:
The vulnerability generalizes. Enron’s undefended hybrid pipeline (P3) produces $\text{RPR}=0.695$ on benign queries—139 of 200 queries leak cross-department content, with a mean of 7.1 leaked items per query. Under adversarial queries, $\text{RPR}=0.615$ (123 of 200 queries leak). The vulnerability is not an artifact of synthetic construction: real corporate email with natural departmental boundaries exhibits the same pivot mechanism.
RPR scales with entity connectivity. The Enron corpus produces lower RPR (0.695 vs. 0.954) and lower mean leakage (7.1 vs. 16.0) than the synthetic corpus. This difference is explained by entity connectivity: the synthetic graph has 40 intentional bridge entities across 4 tenants with curated cross-references, while Enron has 19 naturally shared entities across 5 departments with sparser cross-department email traffic. The Enron graph is $135\times$ larger (376K vs. 2.8K nodes) but its entity mentions are noisier and more fragmented, reducing the density of cross-tenant pivot paths. This confirms that RPR is a function of entity connectivity density, not corpus size.
D1 eliminates all leakage on both corpora. P4 achieves $\text{RPR}=0.0$ on Enron with context size 25 (vs. 56 on the synthetic corpus). The smaller post-D1 context reflects Enron’s higher entity-to-chunk ratio—more entity nodes are filtered, leaving fewer authorized chunk traversals. The defense generalizes without modification: the same per-hop tenant-and-sensitivity check eliminates leakage regardless of corpus size, entity distribution, or graph topology.
PD = 2 persists. All leakage in both corpora occurs at exactly $\text{PD}=2$, confirming that the bipartite pivot (chunk $\to$ entity $\to$ chunk) is a structural invariant of hybrid RAG systems that construct knowledge graphs via entity linking, independent of the underlying document domain.
VIII Mitigations
We propose five defenses that operate at different stages of the hybrid pipeline, forming a defense-in-depth architecture (Table XIV).
TABLE XIV: Defense mechanisms, pipeline integration points, and experimental impact.
Defense Stage Effect
D1: Per-hop authz Post-expansion RPR $\to$ 0.0
D2: Edge allowlist Traversal query Defense in depth
D3: Budget Traversal params Ctx 110 $\to$ 28
D4: Trust weight Post-expansion Low-prov removal
D5: Merge filter Post-merge Final backstop
VIII-A D1: Per-Hop Authorization
Per-hop authorization is an intentionally conservative minimum-viable guardrail: it re-checks access control predicates on every node reached during graph expansion. For each expanded node $v$, the defense evaluates:
$\text{auth}(u,v)=(v.\text{tenant}=t_{u})\wedge(v.\text{sensitivity}\leq\ell_{u})$ (8)
Nodes failing this check are removed from the expansion result before context assembly. In our implementation, D1 operates as a post-expansion filter: the BFS traversal runs unrestricted, and the authorization check is applied to the result set. This is simpler to implement than in-traversal pruning (which would stop expansion at unauthorized nodes) but is less efficient because it expands nodes that will be discarded. A true per-hop pruning implementation would stop expansion at unauthorized nodes, preventing their children from being explored. We chose post-expansion filtering for implementation simplicity; the security properties are identical (both produce $\text{RPR}=0.0$), but in-traversal pruning would further reduce latency and context processing overhead.
D1 is the most effective defense: it reduces RPR from 0.95 to 0.0 across all query types and all attack variants, with <<1ms latency overhead.
Entity tenant semantics.
A subtle but critical design decision concerns entity nodes. In our knowledge graph, entity nodes (extracted via NER) represent shared concepts—people, systems, vendors—that may appear in documents across multiple tenants. These nodes carry tenant = "" (empty string) rather than any specific tenant label, because they are tenant-neutral: the concept “CloudCorp” belongs to no single tenant. Under D1’s authorization check, empty-tenant nodes fail the tenant-match predicate ($v.\text{tenant}\neq t_{u}$) and are filtered from the expansion result. This means D1 removes all entity nodes from the context—not just unauthorized ones.
This entity-level filtering is precisely the mechanism by which D1 eliminates cross-tenant leakage. The 2-hop pivot path (chunk $\to$ entity $\to$ chunk) requires traversal through a shared entity node. By filtering entity nodes, D1 severs this path at hop 1, preventing the traversal from reaching the unauthorized chunk at hop 2. The trade-off is that entity information (which may be useful for answer generation) is excluded from the final context. We quantify this impact in Section IX.
We acknowledge that D1’s entity filtering is functionally equivalent to disabling entity traversal—a valid criticism. However, this is the minimum viable defense, not the optimal one. A finer-grained entity-aware authorization scheme would: (a) label each entity with the set of tenants whose documents mention it (e.g., “CloudCorp” $\to$ {acme, globex}), (b) allow traversal through entities whose tenant set includes $t_{u}$, and (c) apply the authorization check only at chunk nodes (the terminal nodes that carry sensitive text). This design would preserve entity context for within-tenant traversals while still blocking cross-tenant chunk access. We leave its implementation and security analysis to future work, and position D1 as the conservative baseline that demonstrates the existence of the boundary problem.
The critical insight is that existing graph databases already store the metadata needed for this check (tenant labels, sensitivity tiers). The defense does not require new infrastructure—only the discipline to re-check authorization at the graph layer rather than relying solely on the vector prefilter.
VIII-B D1 Metadata Integrity Assumption
D1 assumes that node metadata (tenant, sensitivity) is trustworthy. If an attacker can forge metadata during document injection, D1 can be bypassed. Our metadata stress test (§VII-K) shows that D1 is robust to random mislabeling up to 5%, but targeted metadata forgery is a different threat. Mitigations include: (a) system-assigned metadata during ingestion (the uploader cannot choose tenant or sensitivity labels), (b) cryptographic signing of metadata at ingestion time, and (c) periodic metadata audits that compare node labels against source-of-truth identity systems. In practice, most enterprise document management systems already enforce system-assigned tenant labels, making (a) the natural deployment model.
VIII-C D2: Edge Allowlist
Edge allowlisting restricts graph traversal to pre-approved edge types per query class. The expander’s BFS query includes a relationshipFilter parameter that limits which edge types can be traversed. For example, general queries may traverse MENTIONS, DEPENDS_ON, and BELONGS_TO edges, while RELATED_TO edges (which are the primary vector for bridge attacks) are excluded.
VIII-D D3: Budgeted Traversal
Budgeted traversal enforces hard limits on BFS expansion: maximum hop depth, maximum branching factor per node, and maximum total expanded nodes. Our traversal sweep (§VII-I) shows that $\text{total\_nodes}\leq 25$ eliminates leakage even without D1, while the default budget ($d_{\text{max}}=2$, branching $\leq 10$, total $\leq 50$) reduces context from 56 items (D1 only) to 28 items, removing authorized but irrelevant graph content that adds noise.
VIII-E D4: Trust-Weighted Expansion
Trust-weighted expansion filters expanded nodes by their provenance score. Each node carries a provenance score (0.00.01.01.0) based on its source reliability. The defense applies a minimum threshold (default: 0.6), removing low-provenance content. This is particularly effective against injected attack payloads, which carry provenance scores of 0.3–0.4 in our attack implementations.
VIII-F D5: Merge-Time Policy Filter
The merge-time policy filter performs a final access control check after vector and graph results are combined. This defense acts as a backstop: even if earlier defenses miss an unauthorized item, the merge filter removes any item whose sensitivity exceeds the user’s clearance before the context reaches the LLM.
IX Discussion
IX-A Security-Utility Tradeoff
Our results reveal a favorable security-utility tradeoff. D1 alone achieves complete security ($\text{RPR}=0.0$) while retaining 50% of the graph-expanded context ($110\to 50\text{--}56$ items). The retained items are authorized content that contributes to answer quality. Adding D3–D5 further reduces context to 19–20 items—an 82% reduction from the undefended baseline—but the additional defenses trade context volume for noise reduction rather than security improvement.
Entity over-filtering.
D1’s security guarantee comes at a specific cost: entity nodes are excluded from the final context. In our bipartite construction, the undefended P3 pipeline returns approximately 110 items per query, of which roughly 25–35 are entity nodes. After D1 filtering, these entity nodes are removed entirely, leaving 50–56 chunk-only items. The entity information they carried (names, relationships, organizational context) is lost. For queries that benefit from entity-level knowledge (e.g., “Who works on project Alpha?”), this may reduce answer quality compared to an oracle that retains only authorized entity nodes. We consider this an acceptable trade-off: the entity nodes are exactly the pivot vectors that enable cross-tenant leakage, and their removal is the mechanism by which D1 achieves $\text{RPR}=0.0$. Future work on multi-tenant entity authorization (Section X) could recover entity utility while preserving security.
The practical recommendation is clear: deploy D1 immediately as a minimum viable defense. Add D3 and D4 if context noise is impacting LLM generation quality. D2 and D5 provide defense in depth for compliance-sensitive deployments. Practically, this means systems should treat tenant-neutral entity nodes as privileged pivot infrastructure that must carry explicit authorization semantics (or be excluded) in multi-tenant deployments; leaving them unlabeled creates a structurally inevitable cross-tenant pivot path.
IX-B Why PD = 2 Across Both Corpora
Proposition (Pivot depth in bipartite entity-link graphs). In a bipartite graph constructed from chunk nodes and entity nodes via NER-based entity linking, any entity co-mentioned by chunks from different tenants induces a length-2 path between those chunks (chunk $\to$ entity $\to$ chunk). Therefore, if cross-tenant leakage occurs via shared entities under expansion depth $d\geq 2$, the minimum pivot depth to the first unauthorized chunk is $\text{PD}=2$.
This follows directly from the bipartite construction: the entity-linking step creates a 2-hop path between any two chunks that mention the same entity. Chunk A (hop 0) $\to$ shared entity (hop 1) $\to$ Chunk B (hop 2). Every shared entity is a potential pivot, and the number of pivot opportunities scales with the number of cross-boundary entity mentions.
The Enron corpus confirms this structural invariant: despite a $135\times$ larger graph (376K vs. 2.8K nodes), noisier NER extractions, and real-world email text rather than curated templates, all leakage occurs at exactly $\text{PD}=2$. The synthetic corpus has 40 shared entities; Enron has 19. Both produce the same pivot depth. This confirms that $\text{PD}=2$ is a consequence of the bipartite construction, not of corpus-specific properties.
Knowledge graphs with richer topologies—ontological hierarchies, multi-hop inference chains, entity-to-entity relationships beyond simple co-mention—may exhibit leakage at PD >2>2. Our traversal sweep confirms that depth=1 prevents all leakage in the bipartite construction, but graphs with entity-entity edges would require analysis at deeper traversal depths.
IX-C The Amplification Mechanics
The core insight of this work is not simply “apply ACLs to graph expansion” but rather the identification and characterization of a compound attack surface that arises from combining two individually secure retrieval modalities. Vector retrieval with tenant prefiltering is secure ($\text{RPR}=0.0$). Graph retrieval within a single tenant’s subgraph is secure. But connecting vector retrieval outputs to graph expansion inputs—the pivot boundary—creates leakage that neither modality exhibits alone.
[…]
The defense implication is equally specific: the authorization check must be placed at the boundary, after graph expansion and before context assembly. Authorization at the vector layer alone (which every production system already has) is insufficient. Authorization at the LLM layer (e.g., instruction-following) is unreliable. Only authorization at the graph expansion layer—the pivot boundary itself—addresses the root cause.
IX-D Comparison with Prior Art
GRAGPoison [8] demonstrates that graph structure amplifies small poisoning inputs (98% ASR from relation injection). Our work extends this insight to hybrid architectures: the combination of vector retrieval and graph expansion creates a compound attack surface where authorized vector results seed unauthorized graph traversals.
PoisonedRAG [28] achieves 90% ASR by injecting 5 documents into vector stores. In vector-only RAG, this is contained by prefilters. In hybrid RAG, even a single retrieved document that mentions a shared entity can trigger graph expansion into sensitive neighborhoods, making the effective attack surface orders of magnitude larger.
The Graph RAG Privacy Paradox [10] observes that graph RAG reduces text leakage but increases structural information extraction. Our findings align: the hybrid pipeline’s graph expansion systematically transforms authorized text into unauthorized structural access.
IX-E Implications for Agentic Systems
The pivot attack is especially dangerous in agentic RAG deployments (LangGraph, CrewAI) where graph traversal is performed autonomously by an LLM agent [6]. In these systems, the agent decides traversal depth, edge types, and expansion strategies without human oversight. A poisoned vector seed can manipulate the agent into performing unrestricted graph exploration, and research shows that a single compromised agent can poison 87% of downstream decision-making within 4 hours [6]. Per-hop authorization is even more critical in agentic settings, where there is no human in the loop at the expansion step.
X Limitations
Two corpora. We evaluate on a synthetic enterprise corpus and the Enron email corpus. The synthetic corpus validates the vulnerability under controlled conditions; the Enron corpus confirms generalization to real-world data. However, both use a bipartite chunk–entity graph topology that produces a uniform PD=2 signature. Knowledge graphs with richer entity-to-entity relationships (ontological hierarchies, multi-hop inference chains) may exhibit leakage at deeper pivot depths. We do not evaluate on production enterprise graphs with RBAC/ABAC policies, which may exhibit different connectivity patterns.
Limited generation evaluation. Our generation contamination experiment (§VII-H) evaluates 10 queries per LLM on the synthetic corpus and shows that leaked context does affect generated answers (ECR up to 0.31, FCR up to 0.19). However, the sample size is small and limited to one corpus. A larger-scale generation study across both corpora would strengthen the end-to-end impact claim. We also do not evaluate chain-of-thought reasoning, which might amplify the impact by synthesizing cross-tenant connections.
Two embedding models. We evaluate two open-source models (all-MiniLM-L6-v2, 384d, and all-mpnet-base-v2, 768d) and confirm the vulnerability persists across both (§VII-E). However, we do not test commercial embedding models (e.g., OpenAI text-embedding-3-large) which may produce different similarity distributions. The pivot vulnerability is structural (entity linking, not embedding quality), so we expect model independence to hold broadly.
Entity over-filtering. D1 removes all entity nodes from the expansion result because entities carry empty-string tenant labels. This is the mechanism that eliminates leakage, but it also removes potentially useful entity context. A finer-grained approach would assign entities to the set of tenants whose documents mention them, or introduce a SHARED tenant class with explicit cross-tenant grants. Such designs would recover entity utility at the cost of more complex authorization logic and a larger attack surface.
Adaptive attacker scope. We evaluate three adaptive attacks (A5–A7) including targeted metadata forgery at rates up to 10%, entity manipulation, and query manipulation (§VII-G). All defended pipelines maintain $\text{RPR}=0.0$. However, we do not evaluate attackers who craft high-provenance payloads to bypass D4 specifically, or adversarial query phrasing to manipulate D2’s query classifier. Our metadata forgery assumes the attacker can relabel node properties but cannot modify the graph schema or index structure.
NER and entity linker dependence. The pivot path exists because entity linking creates shared nodes across tenant boundaries. Both corpora use spaCy’s en_core_web_sm NER model, which has known recall limitations for domain-specific entities. On the Enron corpus, this produces noisier entity extractions (dates, monetary amounts, partial names) that create spurious cross-tenant connections—yet RPR is lower (0.695 vs. 0.954), suggesting that NER noise creates fragmented rather than dense pivot paths. A higher-recall linker would extract more shared entities, potentially increasing organic leakage. We do not evaluate how linker quality affects RPR, nor do we test production entity linking systems that perform coreference resolution or cross-document entity merging.
Simplified policy model. Our authorization model uses tenant labels and four sensitivity tiers (PUBLIC through RESTRICTED). Production enterprises typically employ richer access control: role-based (RBAC) and attribute-based (ABAC) policies, group memberships, temporary grants, legal holds, and need-to-know exceptions. D1’s per-hop check generalizes to any predicate that can be evaluated on node metadata, but we do not demonstrate this generality experimentally. The gap between our flat tenant model and real-world RBAC/ABAC hierarchies remains an open integration question.
Graph-only baseline (P2). We omit a graph-only RAG baseline because graph-only retrieval without vector seeding does not involve the pivot boundary under study. This means our results characterize the hybrid-specific vulnerability but do not compare against graph-only retrieval’s independent security properties.
XI Related Work
XI-A Vector RAG Poisoning
PoisonedRAG [28] formalized knowledge corruption attacks on RAG, achieving 90% ASR with 5 injected documents. CorruptRAG [24] demonstrated single-document attacks with higher stealth. CtrlRAG [18] achieved 90% ASR on GPT-4o via black-box feedback optimization. RIPRAG [21] applied reinforcement learning to optimize poisoning without model access. NeuroGenPoisoning [27] targeted specific neurons for ¿90% success. These works study vector-side attacks in isolation; none address what happens when poisoned vector results seed graph expansion.
XI-B GraphRAG Security
GRAGPoison [8] is the closest related work, demonstrating relation-centric poisoning of GraphRAG with 98% ASR. TKPA [20] showed that modifying 0.06% of corpus text drops GraphRAG accuracy by 45 percentage points. RAGCrawler [22] achieved 84.4% knowledge extraction coverage through graph-guided probing. The Graph RAG Privacy Paradox [10] established that graph RAG increases structural leakage while reducing text leakage. Our work extends these graph-side insights to the hybrid setting, where the vector-to-graph transition creates a compound attack surface.
XI-C RAG Privacy and Extraction
The SoK on RAG Privacy [2] systematized all known privacy attack vectors in RAG systems and explicitly noted that hybrid RAG privacy risks remain under-studied. Riddle Me This [12] demonstrated membership inference on RAG systems. Traceback of RAG Poisoning [23] provided forensic methods for identifying responsible documents, but traces only through the vector retrieval path without following graph expansion chains.
XI-D RAG Defenses
RAGuard [3] detects poisoning via retrieval pattern analysis. SeCon-RAG [17] applies semantic consistency filtering. RevPRAG [19] achieves 98% detection via LLM activation analysis. SDAG [5] partitions context into trusted and untrusted segments with block-sparse attention. SD-RAG [1] implements selective disclosure policies. All operate within a single retrieval modality. Our defense suite (D1–D5) is the first designed specifically for the cross-store boundary, with per-hop authorization as the cornerstone mechanism.
XI-E Ethical Considerations
Experiments use two corpora: a synthetic corpus generated by the authors and the Enron email corpus, which is a public-record dataset released during the 2001 FERC investigation and widely used in NLP research [4]. No production enterprise data, user accounts, or confidential documents were used. The attack implementations (A1–A7) target our own evaluation infrastructure and are designed to characterize vulnerabilities, not exploit production systems. Our primary contribution is the defense (D1), which we release alongside the attack code to ensure that the net effect of publication is protective. The released code does not include tools for targeting external systems; all components require a local Neo4j and ChromaDB deployment to operate.
XII Conclusion
Hybrid RAG pipelines that combine vector retrieval with knowledge graph expansion introduce an attack surface at the vector-to-graph boundary that has received limited explicit treatment in prior work. We formalized this threat as Retrieval Pivot Risk (RPR) and demonstrated across two corpora—a synthetic enterprise dataset (1,000 documents, 4 tenants) and the Enron email corpus (50,000 emails, 5 departments)—that undefended hybrid pipelines exhibit $\text{RPR}=0.95$ and $\text{RPR}=0.70$ respectively, compared to $0.0$ for vector-only retrieval. All leakage occurs at 2 hops through the entity-linking pivot in our bipartite graph construction, a structural invariant that holds across both corpora.
Our taxonomy of seven Retrieval Pivot Attacks (A1–A4 non-adaptive, A5–A7 adaptive) shows that small injection budgets (10–20 chunks) can exploit this boundary, and that even without adversarial injection, naturally shared entities create organic pivot paths that leak cross-tenant data. The Enron corpus confirms this on real data: 19 naturally shared entities across departmental boundaries produce RPR == 0.695 on benign queries alone.
The most important finding is practical: per-hop authorization—re-checking tenant and sensitivity labels at each graph expansion step—eliminates all measured retrieval pivot leakage across both corpora, all seven attack variants, and metadata forgery rates up to 10%. This defense is simple, requires no model changes, adds <<1ms latency, and uses metadata already present in graph databases. We recommend it as the minimum viable security control for any hybrid RAG deployment.
Our traversal parameter sweep reveals that the total node budget is the primary leakage-controlling parameter: limiting expansion to $\leq$25 nodes eliminates leakage even without authorization checks, while expansion depth must reach $\geq$2 for the pivot to occur. The full defense stack (D1–D5) reduces context noise by 82% with zero residual leakage.
The cross-dataset comparison reveals that RPR scales with entity connectivity density: the synthetic corpus (40 bridge entities, curated cross-references) produces higher RPR than Enron (19 natural bridges, sparser cross-department traffic). This confirms that the vulnerability is structural but its severity depends on the knowledge graph’s topology—denser entity sharing means more pivot opportunities.
Our work opens several directions for future research: evaluation on production enterprise graphs with richer entity-to-entity topologies (where PD >> 2 may emerge), high-provenance payload crafting to bypass D4, multi-tenant entity authorization (recovering entity context without re-enabling the pivot path), and extension to agentic hybrid RAG where graph traversal is autonomously controlled.
Appendix A Artifact Appendix
A-A Repository
The complete codebase, synthetic data generators, attack implementations, defense suite, evaluation harness, and experimental results are available at: https://github.com/scthornton/hybrid-rag-pivot-attacks
A-B System Requirements
  • Python 3.11+ with dependencies: chromadb, neo4j, spacy, sentence-transformers, pydantic, numpy
  • Neo4j 5.15+ with APOC plugin (for apoc.path.spanningTree)
  • ChromaDB server (latest)
  • spaCy model: en_core_web_sm
A-C Reproduction Steps
Synthetic corpus:
  1. 1.
    Clone repository and install: pip install -e ".[dev]"
  2. 2.
    Start services: docker compose up -d (Neo4j + ChromaDB)
  3. 3.
    Generate corpus: python scripts/make_synth_data.py
  4. 4.
    Generate queries: python scripts/generate_queries.py
  5. 5.
    Build indexes: python scripts/build_indexes.py
  6. 6.
    Run experiments: python scripts/run_experiments.py --bootstrap
  7. 7.
    Run attack experiments: python scripts/run_attack_experiments.py
  8. 8.
    Run sweeps: python scripts/run_sweep_experiments.py --traversal-sweep --mislabel-sweep --connectivity-sweep
Enron corpus:
  1. 1.
    Ingest Enron emails: python scripts/ingest_enron.py
  2. 2.
    Build indexes: python scripts/build_indexes.py --dataset enron
  3. 3.
    Run experiments: python scripts/run_experiments.py --dataset enron
Run tests: pytest tests/ -v (255 tests passing).
A-D Runtime Estimates
Synthetic: Corpus generation and index building complete in $\sim$5 minutes. The full 500-query evaluation across 7 pipeline variants takes $\sim$20 minutes. Attack experiments ($4\times 4$ configurations) require $\sim$10 minutes including graph rebuilds. Sweep experiments total $\sim$45 minutes.
Enron: Ingestion and NER extraction takes $\sim$40 minutes (2.07M entity mentions across 152K chunks). Neo4j graph loading completes in $\sim$2 minutes using batch UNWIND queries. The 200-query evaluation across P1/P3/P4 takes $\sim$10 minutes.
All experiments require <<16 GB RAM and run on CPU-only hardware.
Appendix B Corpus Generator Details
The synthetic corpus is produced by 12 domain-specific document generators (3 per tenant), each instantiating one of three genre templates: architecture description, policy/procedure, and status report. Each generator populates templates from curated entity pools:
  • Engineering: 12 system names, 15 technologies, 6 projects (e.g., auth-service, Kubernetes, Project Alpha).
  • Finance: 10 vendors, 6 accounts, 7 regulations (e.g., Deloitte, SOX, Capital Expenditure 2025).
  • HR: 12 departments, 7 benefits, 10 named employees (e.g., Engineering, 401k matching, Maria Chen).
  • Security: 6 CVEs, 8 tools, 6 frameworks (e.g., CVE-2025-41923, Splunk SIEM, NIST CSF).
Bridge entities span 5 categories: shared vendors (CloudCorp, DataSyncInc, SecureNetLLC), shared infrastructure (k8s-prod-cluster, splunk-siem, auth-service), shared personnel (Maria Chen, James Rodriguez, Aisha Patel), shared compliance (SOC2-audit, PCI-DSS-cert, ISO27001), and shared projects (ProjectNexus, ProjectHorizon, ProjectArcade). Sensitivity tiers are assigned per-document: PUBLIC 40%, INTERNAL 30%, CONFIDENTIAL 20%, RESTRICTED 10%. All generation uses seed 42 for reproducibility.
Appendix C Query Templates
Benign queries (350 total, 5 templates $\times$ 4 domains $\times$ variable entity fills). Example templates:
Engineering: "What is the architecture of {system}?"
Finance:     "Summarize the contract with {vendor}."
HR:          "What are the policies for {department}?"
Security:    "What is the severity of {cve}?"
Adversarial queries (150 total) are stratified across 4 attack types:
A1 (Seed Steering):     "What systems like {system}
     handle {other_tenant_technology}?"
A2 (Entity Anchor):     "Tell me about {bridge_entity}
     and its role in {target_tenant_domain}."
A3 (Neighborhood Flood): "List everything related to
     {bridge_entity}."
A4 (Bridge Node):       "What connections exist between
     {bridge_entity} across departments?"
Template slots are filled from aligned entity pools (identical to those in the corpus generator) to ensure query entities appear in the knowledge graph. Each query carries metadata: user_tenant (always acme_engineering), user_clearance (stratified across PUBLIC, INTERNAL, CONFIDENTIAL), and query_type (benign or adversarial with attack subtype).
Appendix D Defense Implementation
D1 (Per-hop authorization). After BFS expansion via apoc.path.spanningTree, each returned node is checked:
def is_node_authorized(node, user):
    tier = SensitivityTier(node.sensitivity)
    if tier > user.clearance:
        return False
    if node.tenant not in user.allowed_tenants:
        return False
    return True
[…]
D3 (Budgeted traversal). The Cypher query enforces a global node budget via LIMIT \$max_total. Per-hop branching is enforced post-query: expanded nodes are grouped by hop depth, and each group is truncated to max_branching entries.
BFS Cypher query (simplified):
UNWIND \$seed_ids AS seed_id
MATCH (start {node_id: seed_id})
CALL apoc.path.spanningTree(start, {
    maxLevel: \$max_hops,
    limit: \$max_total
}) YIELD path
WITH last(nodes(path)) AS node,
     length(path) AS depth
RETURN node.node_id, node.tenant,
       node.sensitivity, min(depth)
ORDER BY hop_depth LIMIT \$max_total
The length(path) return value feeds the Pivot Depth metric directly.
Appendix E Entity-Aware Authorization
D1 eliminates all entity nodes because they carry empty-string tenant labels. This is effective but coarse: it removes potentially useful entity context (names, types, relations) from the retrieval result. A finer-grained scheme would distinguish traversal authorization from inclusion authorization:
  1. 1.
    Entity tenant-set labeling. During graph construction, label each entity with the set of tenants whose documents mention it: tenant_set = {acme, globex}. Shared entities retain multiple tenant labels.
  2. 2.
    Traversal-through permission. Allow BFS to traverse through an entity node if the user’s tenant is in the entity’s tenant_set. The entity itself may appear in context.
  3. 3.
    Destination check. At each chunk node reached via an entity, re-check the chunk’s tenant and sensitivity against the user’s policy. This preserves D1’s security guarantee at the chunk level while recovering entity utility.
This scheme would increase context size (recovering the entity nodes D1 currently removes) without re-enabling the $\toentity\tounauthorized chunk$ pivot, because the destination check still blocks cross-tenant chunks. The key trade-off is implementation complexity: entity tenant_set must be maintained as documents are added or removed, and the authorization predicate becomes a set-membership check rather than a simple equality.
Appendix F Statistical Methodology
Bootstrap procedure. All confidence intervals use the non-parametric percentile bootstrap (10,000 resamples, seed 42). For binary RPR indicators, the bootstrap distribution is constructed by resampling the $n=500$ per-query binary outcomes $\{0,1\}$ with replacement and computing the mean for each resample. The 95% CI is the $2.5\%,97.5\%]$ percentile interval of the bootstrap distribution.
Sample size justification. With $n=500$ queries, we achieve power >0.80>0.80 for detecting a 5 percentage-point difference in RPR (from 0.0 to 0.05) at $\alpha=0.05$, computed via the exact binomial test. For leakage means, the standard error of the bootstrap is $\text{SE}<0.5$ items, sufficient to distinguish mean differences of $\geq 1$ item.
Multiple comparisons. We test 7 pipeline variants against 2 query types (14 comparisons per metric). Applying a Bonferroni correction ($\alpha_{\text{adj}}=0.05/14=0.0036$), the core findings (P1/P4–P8: RPR =0.0=0.0; P3: RPR >0.90>0.90) remain significant at $p<10^{-6}$.
$\epsilon$ sensitivity. The regularized amplification factor $\text{AF}(\epsilon)$ uses $\epsilon=0.1$. We verify that conclusions are robust across $\epsilon\in\{0.01,0.05,0.1,0.5\}$: $\text{AF}(\epsilon)$ ranges from 1,599 ($\epsilon{=}0.01$) to 32 ($\epsilon{=}0.5$), all confirming $30\times$ amplification.